A projectile is launched horizontally from the top of a vertical cliff with a velocity of 15m/s is observed to strike the ground at a distance of 45m from the foot of the cliff. What is the angle the path of the projectile makes with the ground after the moment of impact


1
Expert's answer
2021-08-22T15:46:51-0400

Range =45m

velocity

vh=15m/sec


Newton's motion 1st law


"v=u+at"

Put u=0,a=g

"v_v=0+9.8\times3=29.43m/sec"

Horizontal velocity

vh=15m/sec

Angle


"\theta=tan^{-1}(\frac{v_v}{v_h})=tan^{-1}(\frac{29.43}{15})=62.99°"